UNIQUE CONTINUATION FOR THE ELASTICITY SYSTEM WITH RESIDUAL STRESS AND ITS APPLICATIONS

GEN NAKAMURA JENN-NAN WANG

Abstract. In this paper we prove the unique continuation property for the elasticity system with small residual stress. The elasticity system becomes anisotropic due to the existence of residual stress. The main technique in the proof is Carleman estimates. Having proved the unique continuation property, we will study the inverse problem of identifying the inclusion or cavity.